(3y−2xy^3)dx+(4x−3x^2y^2)dy=0


1
Expert's answer
2021-06-03T15:24:00-0400

Given


"(3y−2xy^3)dx+(4x−3x^2y^2)dy=0\\"

then


"M(x,y) = 3y−2xy^3\\
N(x,y) =4x−3x^2y^2\\
\frac{\partial M}{\partial y}=3-6xy^2\\
\frac{\partial N}{\partial x}=4-6xy^2\\
\frac{\partial M}{\partial y}\neq\frac{\partial N}{\partial x}\\
\text{Showing that the equation is not exact.}\\"

Let

"x^ay^b"

be the integrating factor such that the given D.E will be exact. Multiplying the given equation through with the I.F gives;


"(3x^ay^{b+1}-2x^{a+1}y^{b+3})dx +(4x^{a+1}y^{b}-3x^{a+2}y^{b+2})dy\\"

Since the equation is now exact, then


"\frac{\partial M}{\partial y}=\frac{\partial N}{\partial x}\\"

where:


"\frac{\partial M}{\partial y} = 3(b+1)x^ay^{b}-2(b+3)x^{a+1}y^{b+2}\\
\frac{\partial N}{\partial x} = 4(a+1)x^ay^{b}-3(a+2)x^{a+1}y^{b+2}\\"



By comparing the RHS of the equation, we get:

"3(b+1) = 4(a+1) \implies 4a-3b=-1 \qquad \cdots eqn(i)\\
-2(b+3)=-3(a+2) \implies b= \frac{3}{2}a \qquad \cdots eqn(ii)"


Solving the two equations simultaneously, we have:


"a=2, b =3"

Thus the IF for the equation to be exact is:


"x^2y^3"

Thus the exact form of the original equation is:


"(3x^2y^4-2x^3y^5)dx+(4x^3y^3-3x^4y^5)dy=0"

Integrating to get the solution to the equation:


"\int (3x^2y^4-2x^3y^6)dx = x^3y^4-\frac{x^4y^6}{2}\\
\int (4x^3y^3-3x^4y^5)dy=x^3y^4-\frac{x^4y^6}{2}"

Therefore, the general solution of the DE is:


"x^3y^4-\frac{x^4y^6}{2}=c\\"